# Vib2Mol: from vibrational spectra to molecular structures—a versatile deep learning model

### v2

@@ 4.4 Alignment between spectrum and molecular structure
where $m$ and $n$ are the number of rows and columns of the probability distribution matrix, respectively. $ilog(p_{i})$andj $ljlog(q_{j})$represent the cross-entropy of spectrum-to-structure, and structure-to-spectrum, respectively. During the testing or inference phase, only the dot product of the features of the to-be-determined spectrum and the molecules in the library needs to be calculated, and the top-k results are taken as the final results (Figure S6).
@@ 4.5 Spectrum-guided molecular generation
4.6 Stochastic Perturbation Beam Search
For de novo generation, the greedy search strategy, which only selects the token with the highest probability as the next character may fall into local optima. To enhance the diversity of the generated results, we adopted a beam search method combined with stochastic perturbation. Specifically, first, the log-probabilities ($logP$) of the current candidate tokens were calculated. Then, Gumbel noise was introduced to perturb these log-probabilities, making the selection process of candidate tokens somewhat stochastic, i.e.:
$logP$ $=log(\frac{e^{\frac{p}{\tau}}}{\sum_{j}{e^{\frac{p_{j}}{\tau}}}})$ (4)
$noise$ $=-log(-log(Uniform(0,1)))$ (5)
$score$ $=logP+\alpha\cdot noise$ (6)
where $p$ is the probability distribution of the current token to be predicted, $\tau$ is the temperature of the random perturbation, a larger $\tau$ can make the probability distribution flatter, thereby increasing the randomness of generation, and $\alpha$ is a weighting factor used to balance the contribution of the log-probability and the current noise. In this method, $\alpha$ is set to 0.1.
4.7 Metrics for conditional generation
4.8 Calculating contributions of different losses
Deep learning models calculate the gradients of the loss function with respect to the weights through chain rule, thereby updating the weights of the entire network. As shown in Figure S5, the CLS token of the spectrum in Vib2Mol is a learnable vector. Therefore, the gradients of each loss function on this vector can represent their corresponding optimization directions, thus quantifying the contribution of each loss function to the weight update. To further compare the differences between the optimization directions guided by each loss function and the total optimization direction, we calculated the cosine similarity between each gradient and the total gradient and normalized it to the range of 0-1, i.e.:
$\nabla L_{i}$ $=\frac{\partial{L_{i}}}{\partial{t}}$ (7)
$CosSim(\nabla L_{i},\nabla L_{total})$ $=\frac{\nabla L_{i}\cdot\nabla L_{total}}{\parallel\nabla L_{i}\parallel\parallel\nabla L_{total}\parallel}$ (8)
$Contribution_{L_{i}}$ $=\frac{CosSim(\nabla L_{i},\nabla L_{total})}{\sum_{j}{CosSim(\nabla L_{j},\nabla L_{total})}}$ (9)
where $L_{i}$ represents each loss fuction, $t$ represents spectral CLS TOKEN, which is a vector.
@@ 4.8 Calculating contributions of different losses
Acknowledgements This work was supported by the National Natural Science Foundation (Grant No: 22227802, 22021001, 22474117, 22272139) of China and the Fundamental Research Funds for the Central Universities (20720220009) and Shanghai Innovation Institute.
Appendix A Reference data
@@
VB-qm9: 133,434 organic small molecules extracted from QM9, composed of C, H, O, N, and F atoms, with the number of heavy atoms less than 10. Each molecule in this subset has only one stable conformation.
VB-zinc15: 50,114 drug molecules extracted from ZINC15[sterling2015zinc], involving a wider range of elements, including C, H, O, N, S, F, Cl, Br, P, and Si, with the number of heavy atoms ranging from 4 to 45. Notably, since the ZINC15 dataset contains many isomers, and VB-zinc15 only ensures the uniqueness of ZINC-IDs, 7,556 molecules in this subset have multiple stable conformations.
@@
VB-geometry: 7,227 organic small molecules extracted from GEOM, each with two stable conformations. We randomly used the spectrum of one conformation as the query input and the other as the reference spectrum, thus constructing a test set for evaluating the model’s spectrum-to-spectrum matching performance.
@@
Appendix B Figures
Figure S4: Normalized cosine values of the gradients with CL, MLM and LM losses relative to the total gradient at different epochs.
Figure S5: Performance of Vib2Mol trained with different mix ratios on unmixed and mixed spectra.
Figure S6: Examples of (A) substitution and (B) Diels-Alder reactions. (C-D) Raman spectra of reactants, product and mixture for each reaction. Compared to the spectrum of the product, the spectrum of the mixture has more characteristic peaks.
Figure S7: Three examples of comparison between theoretical and experimental spectra sampled from NIST.
@@ Appendix B Figures
Appendix C Tables
Table S5: Add caption
VB-geometry VB-qm9 VB-zinc15 VB-mols VB-qm9 VB-zinc15 VB-mols VB-qm9 VB-zinc15 VB-mols
Spectral retrieval Spectrum-structure retrieval Conditional generation De novo generation
(Recall@1 / %) (Recall@1 / %) (molecular accuracy / %) (Recall@1 / %)
Cosine similarity 34.67 / / / / / / / / /
$\rho$ similarity 34.71 / / / / / / / / /
vanilla-CL 76.9 80.94 75.19 78.07 / / / / / /
vanilla-MLM / / / / 97.22 83.06 90.67 / / /
vanilla-LM / / / / / / / 58.65 46.05 54.99
Vib2Mol 76.46 81.51 76.65 78.92 97.91 87.04 92.88 63.31 49.86 59.41
Vib2Mol
(w/o RR) 76.46 81.51 76.65 78.92 97.91 87.04 92.88 60.63 47.68 56.88
Vib2Mol
(w/o RR and SPT) 72.98 77.38 74.29 75.35 97.14 87.92 92.85 62.89 48.34 58.68
Table S6: Performance of 6 methods at 4 tasks during the ablation experiments.
VB-geometry VB-qm9 VB-zinc15 VB-mols VB-qm9 VB-zinc15 VB-mols VB-qm9 VB-zinc15 VB-mols
Spectral retrieval Spectrum-structure retrieval Conditional generation De novo generation
(Recall@1 / %) (Recall@1 / %) (molecular accuracy / %) (Recall@1 / %)
CL-only 76.46 81.51 76.65 78.92 / / / / / /
MLM-only / / / / 97.75 88.74 93.59 / / /
LM-only / / / / / / / 58.65 46.05 54.99
CL+MLM 75.63 78.33 75.53 76.26 98.27 89.13 94.03 / / /
CL+LM 73.39 72.67 70.07 70.66 / / / 61.89 47.91 57.84
CL+MLM+LM 72.98 77.38 74.29 75.35 97.14 87.92 92.85 62.89 48.34 58.68
Table S7: Performance of Vib2Mol in PAHs.
Categories/ Benzene Naphthalene Anthracene Averaged
Recall@1/
Tasks
Substitution site retrieval 99.57 99.75 99.14 99.54
(given substituents)
Substituent generation 99.57 98.72 98.71 98.95
(partially unknown)
Substituent generation 98.28 96.55 95.07 96.64
(both unknown)
De novo generation 95.32 82.44 84.91 86.63
(totally unknown)
@@ Appendix C Tables
Datasets # Training samples # Evaluation samples # Testing samples
qm9 93403 13344 26687
zinc15 38089 5442 10883
geometry 0 0 11318
PAHs 3006 430 860
RXN 10947 1564 3128
Peptide 18168 2596 5191
Peptide-mod 8787 1256 2511
NIST-IR (Exp.) 9055 1046 2043
Table S9: All substituents existing in VB-PAHs.
Name Formula SMILES
Trichloromethyl $CCl_{3}$ C(Cl)(Cl)Cl
Dichloromethyl $CHCl_{2}$ C(Cl)Cl
Chloride $Cl^{-}$ Cl
Chloromethyl $CH_{2}Cl$ CCl
Thiyl (Sulfur) $SH$ S
Trifluoromethyl $CF_{3}$ C(F)(F)F
Difluoromethyl $CHF_{2}$ C(F)F
Fluoride $F^{-}$ F
Fluoromethyl $CH_{2}F$ CF
Hydroxyl (Hydroxy) $OH^{-}$ O
Sulfonic Acid $-OSO_{3}H$ S(=O)(=O)O
Phosphonic Acid $PO(OH)_{2}$ P(=O)(O)O
Nitro $-NO_{2}$ N+[O-]
Carboxylic Acid $-COOH$ C(=O)O
Ketone (Carbonyl) $-CO-$ C=O
Amino $-NH_{2}$ N
Cyano $-CN$ C#N
Methyl $CH_{3}-$ C
Isobutyryl $-C(CH_{3})_{2}O$ C(C)=O
Oxime $-C=NOH$ C(=N)O
Methoxy $OCH_{3}$ OC
Amide $-CONH_{2}$ C(N)=O
Ethyl $CH_{2}CH_{3}$ CC
Propyl $CH_{2}CH_{2}CH_{3}$ CCC
Vinyl $CH_{2}=CH-$ C=C
Ethynyl $C\equiv CH$ C#C

### v3

@@ 4.4 Alignment between spectrum and molecular structure
where $m$ and $n$ are the number of rows and columns of the probability distribution matrix, respectively. $ilog(p_{i})$andj $ljlog(q_{j})$represent the cross-entropy of spectrum-to-structure, and structure-to-spectrum, respectively.
During the testing or inference phase, only the dot product of the features of the to-be-determined spectrum and the molecules in the library needs to be calculated, and the top-k results are taken as the final results (Figure S4).
4.5 Spectrum-structure matching and re-ranking
Building on a large-scale spectrum-structure alignment achieved through contrastive learning, we found that a small number of negative pairs—those with similar spectra or molecular structures—were still difficult to distinguish. To address this, we introduced the spectrum-structure matching. This module is a binary classification task, where the model uses a linear layer to predict whether a spectrum-structure pair is positive (matched) or negative (unmatched) given their multimodal feature. In order to find more informative negatives, we adopt the hard negative mining strategy by BLIP[32], where negatives pairs with higher contrastive similarity in a batch are more likely to be selected to compute the loss. During inference, this module naturally scores the retrieved or generated molecular structures against a to-be-determined spectrum, enabling a deep learning-based re-ranking of candidate molecules.
@@ 4.6 Spectrum-guided molecular generation
4.7 Beam Search
For de novo generation, the greedy search strategy, which only selects the token with the highest probability as the next character may fall into local optima. To enhance the diversity of the generated results, we adopted the beam search strategy, of which the implement is derived from PBSA.
Our method works as follows: at each decoding step, the model calculates the log-probabilities for all possible next tokens. Instead of picking only the most probable token, it keeps track of the top k most probable sequences, where k is the beam size. These sequences are then expanded in the next step, and the process is repeated. The scores for the candidate sequences are accumulated as the sum of their log-probabilities. This parallel exploration of multiple promising paths helps to find better solutions that a greedy approach might miss. To control the generated results, a temperature parameter ($\tau$) is applied to the log-probabilities before the top-k selection. A higher temperature value makes the probability distribution flatter, increasing the randomness and diversity of the selected tokens. Conversely, a lower temperature value sharpens the distribution, leading to a more deterministic search similar to greedy decoding. This allows us to balance between fidelity to the most likely sequence and the exploration of diverse alternatives.
This process can be described by following equations:
$P(w_{t}|W_{<t},X)$ $=\text{softmax}\left(\frac{\text{logits}(w_{t}|W_{<t},X)}{\tau}\right)$
$\text{Score}(W_{T})$ $=\sum_{t=1}^{T}\log P(w_{t}|W_{<t},X)$
where $W_{T}=(w_{1},w_{2},...,w_{T})$ is a candidate sequence of length $T$, $X$ is the input sequence, $W_{<t}=(w_{1},w_{2},...,w_{t-1})$ represents the sequence of tokens generated up to step t1t-1, and $\text{logits}(w_{t}|W_{<t},X)$ are the unnormalized log-probabilities for the next token $w_{t}$.
@@ 4.7 Beam Search
Acknowledgements This work was supported by the National Natural Science Foundation (Grant No: 22227802, 22021001, 22474117 and 22272139) of China and the Fundamental Research Funds for the Central Universities (20720220009 and 20720250005) and Shanghai Innovation Institute.
Appendix A Details about Datasets
@@
VB-QM9: 133,434 organic small molecules extracted from QM9[33], composed of C, H, O, N, and F atoms, with the number of heavy atoms less than 10. Each molecule in this subset has only one stable conformation
VB-ZINC15: 50,114 drug molecules extracted from ZINC15[34], involving a wider range of elements, including C, H, O, N, S, F, Cl, Br, P, and Si, with the number of heavy atoms ranging from 4 to 45. Notably, since the ZINC15 dataset contains many isomers, and VB-zinc15 only ensures the uniqueness of ZINC-IDs, 7,556 molecules in this subset have multiple stable conformations.
@@
VB-geometry: 6835 organic small molecules extracted from GEOM[35], each with two stable conformations. We randomly used the spectrum of one conformation as the query input and the other as the reference spectrum, thus constructing a test set for evaluating the model’s spectrum-to-spectrum matching performance.
SDBS: 2815 organic molecules extracted from Spectral Database for Organic Compounds (SDBS[17]), which contains experimental Raman and infrared spectra simultaneously. This data was collected up to July 1, 2025.
NIST-IR: 12144 experimental infrared spectrum-molecule pairs extracted from NIST Chemistry WebBook[18]. This data was collected up to July 1, 2025.
@@
Appendix B Ablation study of MLM
The MLM model initially achieved a small improvement when we switched from an encoder-only framework (like BERT) to an encoder-decoder architecture. By increasing the training masking ratio from 15% to 45%, we observed a substantial performance leap from 87.91% to 92.46% (Figure S6A). As Figure S6B illustrates, a model trained with a specific masking ratio tends to perform best when tested with a similar masking ratio. For instance, a model trained with 15% masking can recover each token with an accuracy of 99.36% if the input SMILES strings are also 15% masked. However, if the strings are masked by 75%, the accuracy drops considerably to 80.65%. In practical scenarios, such as the prediction of polycyclic aromatic hydrocarbon functional groups discussed in Section 2.5 of main text, the masking rate of input strings can fluctuate significantly and is not fixed. 45% emerged as an optimal hyperparameter because models trained with a 45% SMILES masking rate achieving the highest average token accuracy, demonstrating a balanced performance across both short and long-range conditional generation tasks. Subsequently, data augmentation applied to SMILES strings, and adopting SPT rather than de novo training, yielded slight improvements in performance. Conversely, the introduction of the LM loss led to a noticeable decline in conditional generation metrics. This suggests that during multi-objective optimization, the LM loss became dominant, negatively impacting the performance of the conditional generation task, which is primarily driven by the MLM loss.
Appendix C Details about multi-modal spectral input
The Vib2Mol architecture is designed to handle multimodal spectral data flexibly. As shown in Figure S7, the spectral input is structured as a $B×C×L$ tensor, where: $B$ is the batch size, $C$ is the number of spectral channels (modalities), $L$ is the spectral dimension (the number of sampling points, which is 1024 in this work). We use a 1D convolution with an equal kernel size and stride to convert the input tensor into patch tokens. This convolutional layer’s input channel count is $C$ and its output channel count is $D$ (the feature dimension of each token, which is 768 in our case).
For a single-modality input, $C=1$. For a multimodal input, $C=2$, as the Raman ($S_{r}aman$) and infrared ($S_{I}R$) spectra are concatenated along the channel dimension ($S_{i}n=concat(S_{r}aman,S_{I}R)$). The key reason we can simply adjust the value of $C$ without other architectural changes is that both Raman and infrared spectra share the same x-axis (wavenumber).
Appendix D Figures
Figure S6: (A) Tanimoto similarity distributions for incorrect predictions. (B) Example spectra and molecular structures of failure cases.
Figure S7: t-SNE visualization of spectral and molecular embeddings. (A) shows the overall embeddings, while figures (B-P) show the embeddings for each functional group.
@@ Appendix D Figures
Figure S11: Ablation studies of MLM. (A) The factors affecting the performance of MLM evaluated by molecular accuracy. (B) The relationship between the masking ratio used for training and testing evaluated by token accuracy. Models perform best when the training and testing masking ratios are similar. This highlights that a 45% masking ratio is an optimal hyperparameter, as models trained with this ratio demonstrate robust performance across a wide range of testing conditions.
Figure S12: Diagram of processing multimodal spectral input.
Appendix E Tables
Table S3: Benchmarking on QM9S
extra info Spectrum-structure retrieval de novo
generation
Raman IR Raman IR
VibraCLIP None 93.68 92.06 - -
PBSA chemical formula - - 78.77 69.21
IR2Mol chemical formula - - 80.55 73.67
SMEN conformation 97.89 97.04 65.78 51.54
Vib2Mol None 97.18 95.00 89.02 82.68
Vib2Mol chemical formula 98.11 96.63 90.91 86.77
Table S4: Benchmarking on VB-Mols
extra info Spectrum-structure de novo
retrieval generation
Raman IR Raman IR
VibraCLIP None 82.37 84.05 - -
PBSA chemical formula - - 69.37 61.46
IR2Mol chemical formula - - 69.88 63.82
SMEN conformation 95.43 94.02 53.08 42.25
Vib2Mol None 93.20 91.38 82.59 78.59
Vib2Mol chemical formula 94.66 93.38 86.74 83.69
Table S5: Benchmarking on VB-GEOM
extra info Spectrum-spectrum de novo
retrieval generation
Raman IR Raman IR
Cosine similarity None 36.62 34.28 - -
Pearson correlation coefficient None 35.86 33.39 - -
VibraCLIP None 74.96 64.43 - -
PBSA chemical formula - - 50.9 47.32
IR2Mol chemical formula - - 0.10 0.07
SMEN None 41.80 49.01 0.00 0.00
Vib2Mol None 77.54 75.33 60.23 56.14
Vib2Mol chemical formula 77.54 75.33 66.86 63.07
Table S6: Benchmarking on SDBS
extra info Spectrum-structure de novo
retrieval generation
Raman IR Raman IR
VibraCLIP None 50.36 43.57 - -
PBSA chemical formula - - 41.29 30.68
IR2Mol chemical formula - - 43.62 41.84
Vib2Mol None 78.01 71.28 24.11 24.82
Vib2Mol chemical formula 90.43 86.17 56.03 52.84
Table S7: Benchmarking on NIST-IR
extra info Spectrum-structure de novo
retrieval generation
VibraCLIP None 45.15 -
PBSA chemical formula - 41.56
IR2Mol chemical formula - 47.57
Vib2Mol None 70.78 34.07
Vib2Mol chemical formula 83.54 55.14
@@ Appendix E Tables
Datasets Data source Modality Training size Evaluating size Testing size
QM9S Theoretical IR+Raman 110992 5842 12982
VB-QM9 Theoretical IR+Raman 93403 13344 26687
VB-ZINC15 Theoretical IR+Raman 38089 5442 10883
VB-mols Theoretical IR+Raman 131492 18786 37570
VB-GEOM Theoretical IR+Raman 0 0 6835*2
SDBS Experimental IR+Raman 2393 140 282
NIST-IR Experimental IR 10327 602 1215
PAHs Theoretical Raman 3006 430 860
RXN Theoretical Raman 10947 1564 3128
Peptide Theoretical Raman 18168 2596 5191
Peptide-mod Theoretical Raman 8787 1256 2511